Due to a car accident you receive a settlement from insurance of $50 a month for the next ten years. What is that worth to you today, assuming a 3% inflation rate?

Show all work.
Answer #1

Monthly Insurance Payout = $ 50 and Tenure of Payouts = 10 years or 120 months

Further, the inflation rate of 3 % per annum functions as the discount rate in this case as it is this very rate which reduces the current value of the future $ 50 payouts.

Monthly Interest Rate = (3/12) = 0.25 %

Therefore, present worth of payouts = 50 x (1/0.0025) x [1-{1/(1.0025)^(120)}] = $ 5178.088